# Can Large Reasoning Models Self-Train?

### v1

@@ Abstract
Scaling the performance of large language models (LLMs) increasingly depends on methods that reduce reliance on human supervision. Reinforcement learning from automated verification offers an alternative, but it incurs scalability limitations due to dependency upon human-designed verifiers. Self-training, where the model’s own judgment provides the supervisory signal, presents a compelling direction. We propose an online self-training reinforcement learning algorithm that leverages the model’s self-consistency to infer correctness signals and train without any ground-truth supervision. We apply the algorithm to challenging mathematical reasoning tasks and show that it quickly reaches performance levels rivaling reinforcement-learning methods trained explicitly on gold-standard answers. Additionally, we analyze inherent limitations of the algorithm, highlighting how the self-generated proxy reward initially correlated with correctness can incentivize reward hacking, where confidently incorrect outputs are favored. Our results illustrate how self-supervised improvement can achieve significant performance gains without external labels, while also revealing its fundamental challenges.
@@ 1 Introduction
Pre-training large-scale neural networks on human-curated corpora (e.g., web-scale texts and carefully curated datasets) has endowed language models with broad general-purpose capabilities (Brown et al., 2020; Rae et al., 2022). However, the availability of human-generated data for pre-training is becoming a limiting factor, particularly as computational resources continue to scale rapidly (Hoffmann et al., 2022; Sevilla et al., 2022). Post-training methods, such as supervised finetuning, offer a way to rapidly improve model capabilities via a much smaller set of highly curated human-generated examples, but such reliance on high quality human-generated examples offers limited scalability (Ouyang et al., 2022; Bai et al., 2022).
In order to scale model capabilities further, reinforcement learning from human feedback can be used. This leverages human preference about model generated outputs, a type of feedback that is much easier to obtain (Ouyang et al., 2022; Kaufmann et al., 2024). Reinforcement learning with verifiable rewards offers a scalable alternative to human feedback by using automatic correctness checks as supervisory signals, an approach that has shown considerable success in reasoning and agentic tasks (DeepSeek-AI et al., 2025; OpenAI et al., 2024). However, this approach often depends on human-designed verifiers, which reintroduce the very scalability issues it seeks to overcome.
Achieving superhuman performance will ultimately require models that train in settings where even humans may not know the correct solutions, and are thus unable to provide feedback on model outputs. In such scenario, models must self-improve—that is, they must be able to assess the correctness of their own outputs and use their own judgment as a feedback signal to refine future generations (Zelikman et al., 2022; Song et al., 2025; Huang et al., 2025). If this process can be sustained iteratively, models have the potential to enhance their performance continually. A particularly promising opportunity for such self-improvement arises in contexts characterized by a positive generation-verification gap (Song et al., 2025), where generating correct solutions is challenging, but verifying their correctness is relatively easy. This gap naturally arises in domains such as mathematics, games, and goal-based agentic tasks, where solutions often require complex reasoning or exploration, yet their correctness can be verified more easily.
@@
In this work, we investigate self-improvement methods for large language models aimed at enhancing their mathematical reasoning abilities without reliance on ground-truth annotations or explicit rule-based verifiers. Specifically, we leverage an intrinsic verification criterion, namely model self-consistency (Wang et al., 2023a), to produce supervisory rewards, capitalizing on the observation that consistency among model-generated answers correlates positively with their correctness. Conceptually similar approaches, relying either directly or indirectly on model-generated self-consistency signals, have been explored previously (e.g., Huang et al. (2023); Prasad et al. (2024)) in mathematical domains. Typically, such approaches have been confined to one or few rounds of improvement. In contrast, the key contribution of this work is to demonstrate that this intrinsic supervisory signal can continuously drive model improvement within a reinforcement learning (RL) framework (Sutton et al., 1998), iteratively enhancing performance without access to external labels.
An appealing feature of our proposed approach is its reliance solely on unlabeled data. Because our method does not depend on external labels, it can also be naturally applied in a “test-time training” fashion. Such flexibility allows models to iteratively boost their performance at inference time. Concurrently to our work, test-time training has also recently been investigated by Zuo et al. (2025).
While the prospect of continual model self-improvement is appealing, it is important to recognize a significant limitation: the model’s self-generated reward serves only as a proxy for underlying correctness. Such proxy rewards may give rise to the phenomenon of reward hacking, whereby the model learns to maximize its self-assigned reward by generating increasingly consistent but potentially incorrect answers. This unintended behavior can reduce or even reverse the initial beneficial correlation between confidence and correctness, ultimately constraining sustained improvement. Understanding the precise dynamics of this process is thus essential for further advancement.
In summary, our contributions are fourfold:
  1. (1)
    Motivated by prior works based on consistency based self-improvement, we introduce a simple yet effective self-training reinforcement learning methodology, Self-Rewarded Training (SRT), that uses consistency across multiple model-generated solutions to estimate correctness during RL training, providing self-supervision signals without labeled data.
  2. (2)
    We empirically demonstrate that at early training stages, SRT achieves performance rivaling standard reinforcement-learning methods trained explicitly on gold-standard answers.
  3. (3)
    We analyze the limitations associated with self-generated rewards, revealing how the model’s reward function, initially correlated with correctness, can degrade to reflect confidence rather than true accuracy, leading to the problem of reward hacking.
  4. (4)
    We propose strategies to mitigate reward hacking, laying the groundwork for effective future approaches to sustaining continual model improvement.
@@ 2 Preliminaries
Let $\pi_{\theta}$ denote a language model parameterized by parameters $\theta$. Given a prompt $x$, the model produces a response $y=(y^{1},y^{2},\dots)$ auto-regressively. Formally, each token in the response sequence is generated according to the conditional probability:
@@
Typically, one optimizes the pass rate defined as the average accuracy of the model $\pi_{\theta}$ across a distribution of prompts $\mathcal{X}$, namely
@@
Taking the gradient of the objective function (3) with respect to $\theta$ and employing a baseline (for variance reduction) (Sutton et al., 1998) leads to the well-known policy gradient formulation:
@@
Here $\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot\mid x)}[r(y^{\prime})]$ is the pass rate for prompt $x$. In practice, the policy gradient in  equation 4 is estimated through Monte Carlo samples, yielding the classical REINFORCE algorithm (Williams, 1992). Recent works have modified this base policy gradient formulation to improve its stability, efficiency, and practicality, resulting in advanced methods such as REINFORCE++ (Hu et al., 2025), GRPO (Shao et al., 2024), PPO (Schulman et al., 2017), and RLOO (Ahmadian et al., 2024a).
Implicit to our method is the notion of a generation-verification gap (Song et al., 2025), where generating correct solutions is hard, but verifying them is easy. Due to space constraints, we present the definition in Appendix A.
@@ 3 Self-Rewarded Training
In the absence of external supervision, language models must rely on themselves to produce supervisory signals. Intuitively, if the model can identify a higher-quality answer among several of its generations, this identified improvement can serve as a training signal. Such situations naturally arise in problems characterized by a positive generation-verification gap, including those in mathematics, logical reasoning, and code generation tasks (Li et al., 2022; Brown et al., 2025; Yu et al., 2024).
Self-supervision via majority voting.
A simple yet effective method for leveraging this gap is majority voting, a technique empirically shown to achieve higher accuracy compared to individual model generations (Wang et al., 2023a). In our setting, majority voting involves: (1) sampling multiple answers per prompt, (2) grouping answers according to their parsed final solutions, and (3) estimating the ground truth answer with the most common solution (the mode). Prior works have demonstrated empirically that fine-tuning models directly on these majority-voted labels can boost performance (Huang et al., 2023; Prasad et al., 2024). However, these methods typically employ a single round of improvement guided by majority decisions, with performance bounded by the offline nature of the training procedure. Motivated by recent impressive performance improvements obtained through reinforcement learning in reasoning domains (OpenAI et al., 2024; DeepSeek-AI et al., 2025), we explore whether this self-supervision approach can be effectively extended and iteratively sustained within a reinforcement-learning loop.
Self-Rewarded Training (SRT).
Motivated by these considerations, we propose framing self-improvement as a reinforcement learning problem, where labels are dynamically derived from the model’s evolving majority votes. At a high level, each RL iteration proceeds as follows: (1) Sample a mini-batch of prompts and generate $n$ answers per prompt using the current model, (2) Determine pseudo-labels by identifying the majority-voted answer for each prompt, (3) Use agreement with the majority-voted answer $y_{\text{majority}}$ as an intrinsic binary reward:
$r(y)=\mathbf{1}[\text{answer}(y)=y_{\text{majority}}],$ (6)
@@
             Generate $n$ solutions $y^{(1)},...,y^{(n)}\sim\pi_{\theta}(\cdot|x)$
@@
(4) Perform a single RL update step on this mini-batch using the derived reward function $r(\cdot)$.
The complete approach is concisely detailed in Algorithm 1. Concretely, our method prescribes a specific form of the reward function using model self-consistency and it is thus compatible with all the common RL training algorithms such as PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024a), REINFORCE(Williams, 1992) and REINFORCE$+$ (Hu et al., 2025). Furthermore, as the number of generations per prompt typically falls in the range 16-64 (Yu et al., 2025), SRT incurs no additional compute cost compared to the label-based version of these algorithms.
@@
As long as majority voting leads to a positive generation-verification gap at each RL iteration, we expect iterative self-rewarding to provide a useful supervisory signal. We describe our observations about the empirical performance of SRT in the following section.
4 Experiments
We design experiments to systematically investigate the strengths and limitations of the SRT algorithm. More precisely, we aim to address the following questions: (1) How effective is SRT compared to standard reinforcement learning methods trained explicitly on ground-truth labels, and how well does this approach generalize to unseen problems? (2) Can self-improvement be sustained iteratively, leading to continuous performance enhancement, or is improvement ultimately bounded? (3) What underlying factors influence the effectiveness of self-improvement? (4) Since our approach does not require labeled data, it is applicable at both training and inference phases; thus, what is the effectiveness of SRT when used to increase performance at the test time? We describe our experimental setup first, and then report our observations.
Experimental setup.
We use a Qwen2.5-Math-7B (Yang et al., 2024) model for all our experiments. Since our algorithm is independent of the choice of the underlying policy gradient algorithm, we choose RLOO (Ahmadian et al., 2024b; Kool et al., 2019) for its simplicity. We implement our algorithm on top of the verl [^1] https://github.com/volcengine/verl framework (Sheng et al., 2024). Unless otherwise mentioned, for every RL iteration, we sample 16 prompts, and for each prompt, we generate 32 rollouts from the policy to compute the majority voting label and subsequently use all such rollouts for training. Our KL loss coefficient is 0.001 unless mentioned otherwise. We use the AdamW (Loshchilov & Hutter, 2019) optimizer with $10^{-6}$ learning rate and 0.01 weight decay for all training runs, and usually do not use any learning rate decay. In practice, not all rollouts for a given prompt may contain a parsable solution. In such cases, we perform majority voting only among the rollouts from which a solution can be extracted. If no such rollouts exist, we discard the prompt and all its rollouts from the loss calculation.
Dataset.
For most of the experiments, we use three training datasets: (1) DAPO (Yu et al., 2025) which contains 17398 math problems, primarily sourced from AoPS  [^2] https://artofproblemsolving.com/, a competitive math training platform, (2) MATH500 (Lightman et al., 2023), which contains 12,000 prompts for training and 500 prompts for validation from the original MATH dataset Hendrycks et al. (2021), (3) AIME 1983-2024 (Veeraboina, 2023), where we filter out the AIME-2024 prompts for validation and use the other 919 prompts for training.
4.1 Self-Training via Majority Voting
Figure 2: (Self-Training Performance) Average over 3 heldout test dataset performance of SRT vs training with ground truth reward. Training for longer with SRT can lead to performance collapse, but at the peak, it leads to around 100% performance improvement over the base model. Earlier in the training, it leads to similar improvement as regular RLOO training with ground truth.
We first investigate question (1): How effective is self-improvement compared to standard RL methods trained explicitly on ground-truth labels, and how well does it generalize to heldout datasets? To address this, we train the base model on each of the training datasets (MATH, DAPO, AIME-1983-2023) and evaluate it on out-of-distribution settings using the following heldout test sets: AIME 24 (Jia, 2024), AIME 25 (Lin, 2025), and AMC (He, 2024). We aggregate performance metrics by averaging model accuracy (pass@1 rate, calculated using 32 rollouts per prompt) across these three distinct datasets to reduce variance. The results are presented in Figure 2; additional results, including the comparison with the baselines, can be found in Appendix C.
Surprisingly, we find that on the MATH and AIME training datasets, the self-supervised SRT method achieves results comparable to reinforcement learning on ground truth labels, despite having no access to ground-truth signals. Notably, the pass@1 scores reported in Figure 2 are evaluated on held-out datasets, demonstrating that our self-training procedure generalizes robustly beyond its training distribution.
The results observed for DAPO are more intricate, however. Specifically, when trained on DAPO, we observe that the performance of the SRT algorithm on the test set initially rises at a rate comparable to regular RL trained on ground-truth answers. Yet, approximately around 400-600 training steps, SRT reaches peak performance and subsequently starts to deteriorate. Regular RL training with ground truth labels, on the other hand, continues improving gradually beyond this point. We revisit and analyze this decline in performance in depth in the subsequent section (Section 4.1.1).
Overall, our findings reveal a striking and somewhat unexpected trend: even without any labeled examples, the performance curve of SRT closely tracks that of RL on gold answers during the initial stages of training. Within statistical uncertainties, we observe that SRT essentially matches the peak test-time pass@1 scores of labeled RL methods when training on MATH and AIME’83-AIME’23 datasets. On the challenging DAPO dataset, SRT is still able to achieve a significant fraction (75%) of the final performance exhibited by RL. Moreover, SRT’s peak performance exhibits a relative improvement of around 100% over the base model on all 3 training datasets.
Importantly, the observed performance collapse can be addressed by early stopping the training. Similar to supervised learning or standard RL training, a small validation set with ground-truth labels can be employed to reliably identify the optimal stopping point and retain the peak performance of SRT. We discuss avoiding model collapse in more depth in Section 5 and Appendix D.
4.1.1 What Happens After SRT Peaks in Performance?

### v2

@@ Abstract
Recent successes of reinforcement learning (RL) in training large reasoning models motivate the question of whether self-training — the process where a model learns from its own judgments — can be sustained within RL. In this work, we study this question using majority voting as a simple self-feedback mechanism. On a comprehensive set of experiments on both synthetic and real reasoning tasks, we find that this basic approach improves not only the model’s reasoning performance, but also its capability of generating better quality feedback for the next RL iteration, driving further model improvement. Yet our analysis also reveals a critical limitation of such a self-training paradigm — prolonged RL with self-reward leads to reward hacking where models learn to maximize training (pseudo-)reward, resulting in sudden and complete performance collapse. Together, these results highlight feedback design as the central challenge and call for future research on mechanisms to enable prolonged self-improvement.
@@ 1 Introduction
Pre-training on human-curated corpora has endowed language models with broad general-purpose capabilities (Brown et al., 2020; Rae et al., 2022), but the supply of such data is becoming a bottleneck as compute scales rapidly (Hoffmann et al., 2022; Sevilla et al., 2022). Reinforcement learning (RL) (Sutton et al., 1998) with verifiable rewards (RLVR) addresses this limitation by using automatic correctness checks, and has already shown success in reasoning and agentic tasks (DeepSeek-AI et al., 2025; OpenAI et al., 2024). Yet, achieving super-intelligence inevitably requires moving beyond domains where humans can provide ground-truth solutions. In these domains, an appealing approach is to have models self-improve— judging the correctness of their own outputs and using this signal to refine future generations (Zelikman et al., 2022; Song et al., 2025b; Huang et al., 2025). If sustained iteratively, and if teacher models can also self-improve, this process could enable continual progress without human supervision — which is crucial step for AI based automatic scientific discovery.
Prior work on self-improvement has mainly relied on supervised fine-tuning (SFT) (Zelikman et al., 2022; Huang et al., 2023) or direct preference optimization (DPO) (Rafailov et al., 2024; Prasad et al., 2024), where the self-labeling rule is updated only a handful of times (e.g., 1–10 rounds) and is generally kept fixed within a training round. These studies show that self-improvement can be effective, but leave open whether it can be sustained over longer horizons. Moreover, they are fundamentally bounded by the verification capabilities of the teacher model used to obtain training supervision, which is kept fixed for extended periods of training. In contrast, RL continuously updates the model on freshly generated (“on-policy”) data (Xu et al., 2024; Tajwar et al., 2024; Lanchantin et al., 2025), a property that has been critical to its success in training reasoning models with verifiable rewards. This success raises a natural question: can self-improvement leverage the same continuous-update paradigm? To study this question, we investigate the setting in which the feedback signal is updated at every gradient step—fundamentally altering the dynamics of self-improvement compared to earlier work.
In this RL-based setup, the choice of feedback mechanism is critical. If feedback is always correct (e.g., perfectly verifying mathematical solutions), the procedure reduces to standard RL with ground-truth supervision. However, in practice, self-feedback is imperfect, and its design determines whether self-training is effective or not. As a first step, we study the simplest possible signal: majority voteWang et al. (2023a) has empirically demonstrated that majority vote tends to have higher accuracy compared to individual generations. Here, we cast the majority vote mechanism as a reward function — granting positive reward to model outputs that match the most common answer. Previous works (Huang et al., 2023; Prasad et al., 2024) have employed majority voting mainly as a mechanism to extract better quality generations from a fixed teacher policy to then distill it into the student model. Zuo et al. (2025) concurrently examined a similar RL procedure with majority voting, but in a different setting, as their focus is on training and testing on the same set of prompts. In contrast, our aim is to use this simple pseudo-label generation mechanism to investigate the validity of RL powered self-training frameworks.
Our key contributions are threefold:
  1. (1)
    Our comprehensive set of experiments demonstrate that this simple mechanism yields measurable gains over the base model on key reasoning metrics such as maj@k and avg@k success rates (Section 4.1). Remarkably, we observe clear improvement in the label generating policy after each gradient step, and this translates to gains over employing labels from a fixed teacher. Moreover, self-training achieves comparable performance to RLVR on 4 different base models.
  2. (2)
    In synthetic tasks where one can control the difficulty of the training dataset, we observe that a simple curriculum-based self-training approach can enable the model to keep climbing on progressively harder tasks without ground-truth labels (Section 4.2). Overall, this signals a promising path towards self-improvement of LLMs.
  3. (3)
    However, prolonged training with this framework consistently teaches models to ignore the prompt entirely and output the same template final answer, which maximizes training reward but leads to complete collapse on test datasets. To our knowledge, we are the first to demonstrate this failure mode of self-improving LLMs by employing self-consistency based training signal. In this work, we analyze the dynamics of model collapse in depth and trace them to the self-reinforcing nature of imperfect feedback.
Overall, these findings identify feedback design as the key challenge that future research should address to sustain self-improvement.
@@ 2 Preliminaries
Let $\pi_{\theta}$ denote a language model parameterized by $\theta$. Given a prompt $x$, the model produces a response $y=(y^{1},y^{2},\dots)$ auto-regressively. Formally, each token in the response sequence is generated according to the conditional probability:
@@
Typically, one optimizes the expected pass rate, defined as the average accuracy across a distribution of prompts $\mathcal{X}$:
@@
Taking the gradient of the objective function (3) with respect to $\theta$ and employing a baseline for variance reduction (Sutton et al., 1998) leads to the well-known policy gradient formulation:
@@
Here $\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot\mid x)}[r(y^{\prime})]$ is the average pass rate for prompt $x$. In practice, the policy gradient in  equation 4 is estimated through Monte Carlo samples, yielding the classical REINFORCE algorithm (Williams, 1992). Recent works have modified this base policy gradient formulation to improve its stability, efficiency, and practicality, resulting in advanced methods such as REINFORCE++ (Hu et al., 2025), GRPO (Shao et al., 2024), PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024), Dr. GRPO (Liu et al., 2025), GSPO (Zheng et al., 2025), etc. Implicit to the training method used in our work is the notion of a generation-verification gap (Song et al., 2025b), where generating correct solutions is hard, but verifying them is easy. We present its definition in Appendix A.
@@ 3 Self-Rewarded Training
      Generate $n$ solutions $y^{(1)},...,y^{(n)}\sim\pi_{\text{label}}(\cdot|x)$
@@
Our objective in this work is to investigate whether reliable training supervision for language models can be generated without external labels. The typical practice for online RL involves generating multiple responses to a prompt, then assigning high or low rewards to each generation according to a ground-truth verifier. In the absence of such a verifier, one might develop a mechanism to derive proxy labels. This mechanism then provides a simple recipe for framing self-improvement as an RL problem. At a high level, each iteration proceeds as follows: (1) Sample a mini-batch of prompts, (2) Determine pseudo-labels, $y_{\text{pseudo}}$, using the mechanism for each prompt, (3) Generate $n$ responses per prompt and use agreement with the derived psuedo-labels as an intrinsic binary reward:
$r(y)=\mathbf{1}[\text{answer}(y)=y_{\text{pseudo}}],$ (6)
@@
and then (4) Perform a single RL update step on this mini-batch using the reward function $r(\cdot)$.
Self-supervision via majority voting.
Among several possible choices for determining reasonably accurate pseudo-labels, we explore the simplest one in this work, majority voting. Majority voting has been empirically demonstrated to have higher accuracy compared to individual model generations (Wang et al., 2023a) and is thus a suitable choice to exploit an LLM’s inherent generation-verification gap (see Appendix Figure 9, where majority voting consistently outperforms the model’s average capabilities across 3 test datasets and therefore can act as a verifier). In our setting, models typically produce a step-by-step chain of thought followed by a final answer (in the case of regular RL training, this final answer is extracted and matched with the ground truth to produce rewards), so one can group all responses by their final answer to determine the majority vote. Concretely, assume we want to generate pseudo-labels using policy $\pi_{\text{label}}$ (which can be any reasonable policy). This procedure then involves: (1) sampling multiple answers per prompt using policy $\pi_{\text{label}}$, (2) grouping answers according to their parsed final solutions, (3) estimating the ground truth answer with the most common solution.
Self-Rewarded Training (SRT).
The general procedure is described in Algorithm 1, which henceforth shall be called Self-Rewarded Training (SRT) in this paper. Since the method prescribes a specific form of the reward function using model self-consistency, it is compatible with all the common RL training algorithms such as PPO, RLOO, REINFORCE, GRPO, etc. We study the quality of generated labels during training by controlling $\pi_{\text{label}}$: setting $\pi_{\text{label}}$ to be the base model recovers our familar setting of learning the majority voting decisions of a fixed model (while still using the current policy’s rollouts for RL training), and setting $\pi_{\text{label}}$ to be the current policy $\pi_{\theta}$ after each gradient step allows us to study whether the quality of the learning signal can be concurrently improved during RL training. In the case where we use the current policy $\pi_{\theta}$ to generate our pseudo-labels, we can reuse them for performing the RL gradient step as well. As the number of generations per prompt typically falls in the range 16-64 (Yu et al., 2025), this variant of SRT incurs no additional compute cost compared to the versions of these algorithms employing ground truth labels. In our work, whether the final answer is correctly formatted and parseable is used to filter responses, but given more compute, one can theoretically employ more sophisticated systems like LLM-as-a-judge (Zheng et al., 2023; Gu et al., 2025) or generative verifiers (Zhang et al., 2025a) to further improve the quality of the training signal. We leave these for future work.
@@
As long as majority voting leads to a positive generation-verification gap at each RL iteration, we expect iterative self-rewarding to provide a useful supervisory signal. We describe our empirical observations in the following section.
4 Experiments and Analysis
Figure 2: (SRT improves both performance and quality of generated labels during training.) We investigate self-training under controlled settings on synthetic reasoning tasks from Reasoning Gym. Remarkably, SRT improves not only the mean accuracy, but the majority voting accuracy as well, which is the source of our training supervision. Improvement in the quality of training signal drives further improvement in performance, as SRT outperforms its variant employing the majority votes from a fixed teacher as proxy labels.
In this section, we present the results of our empirical study. Our primary aim is to answer the two following research questions: (1) Can SRT improve an LLMs’ reasoning abilities beyond the base model’s capabilities, in terms of both the quality of training reward signal and performance on downstream tasks? Specifically, since simply performing majority voting on the base model can improve performance compared to individual generations, can we improve the quality of the majority votes (which acts as the verifier for SRT) in addition to pass@1 performance compared to the base model via SRT? (2) If SRT can go beyond the base model, can this improvement be sustained indefinitely? We systematically design experiments to answer these questions below.
4.1 Can SRT go beyond the base model’s capabilities?
There are two potential axes of improvements over the base model for SRT that can be studied: (1) the improvement in accuracy over held-out prompts, and (2) the improvement in the quality of generated labels themselves during the training procedure. Unlike previous works (Huang et al., 2023; Prasad et al., 2024) that distill the majority voting decision of a fixed policy (typically the base model) into the current model, we want to study whether the quality of the majority votes of the evolving policy improves as a result of self-improvement.
Figure 3: (Evaluating SRT on real-world math problems.) Comparison between SRT and RL with ground truth across different base models and training datasets. Following (Oertell et al., 2024), all models are trained using RLOO (for experiments with GRPO, see Figure 18) and tested using average pass@1 accuracy on MATH-500. SRT achieves comparable performance to that of ground-truth training across different base models. For training curves using more combinations of (train, test) dataset pairs, refer to Appendix C.1 and C.2.
Figure 4: (Majority@32 accuracy comparison between SRT and RL with ground truth) We compare the majority@32 accuracy, as opposed to average accuracy shown in Figure 3. Note that for Llama-3.1-8B-Instruct, we use the official model card evaluation temperature of 0, hence majority@32 is the same as average@32 accuracy. SRT shows improvement in the quality of the majority votes themselves, which distinguishes our algorithm from that of learning from a fixed teacher’s majority votes.
Experiments on synthetic reasoning tasks.
Complex reasoning tasks like math require domain-specific pre-training/midtraining for online RL to be effective (Wu et al., 2025a; Gandhi et al., 2025); it is also more difficult to control the difficulty of the individual tasks that the model sees during training. Therefore, we first study this question using synthetic reasoning tasks from REASONING GYM (Stojanovski et al., 2025), a collection of over 100 reasoning environments for reinforcement learning with verifiable rewards. These tasks can also be generated with adjustable difficulty, making it a suitable test-bed for our work. Concretely, we use 3 tasks from Reasoning Gym: (1) Family Relationships, a logic puzzle involving a group of individuals connected via different relationships, and the model has to reason about the relationship between two individuals within this group, (2) Bitwise Arithmetic, a task for testing models’ understanding of Bitwise Arithmetic operations, and (3) Knights & Knaves, a logic puzzle involving characters who always either tell the truth (knights) or always lie (knaves), and the challenge is to deduce who is who based on their statements. Appendix K shows example prompts from each of these tasks. We use a Qwen-3-4B-Base (Yang et al., 2025) model for all our reasoning gym experiments.
Since our goal is to study whether SRT can improve performance on top of a reasonably strong base model, following the setting of Lee et al. (2025), we first train the base model with GRPO using ground truth labels on the easiest difficulty setting. This is used to teach the base model proper formatting rules and how to solve the basic task before we train using SRT on the next level of difficulty without ground-truth labels. The detailed training settings can be found in Appendix B.3.
Results.
Figure 2 shows our main results: SRT using the current policy as the label generator improves both avg@16 and majority vote@16 accuracy on all 3 reasoning gym tasks. Since we derive our training signal from the majority votes of the current policy evolving with every RL step — this demonstrates that self-improvement using SRT can progressively improve the quality of the pseudo-labels as well. Note that this would not be possible in prior works (Prasad et al., 2024; Huang et al., 2023) which distills the majority votes of a fixed teacher policy throughout one round of training. We expect the improvement in the evolving teacher policy to result in further performance gain. To validate this, we compare SRT with a variant of the same algorithm where we use the majority votes of the fixed starting policy instead of the evolving current policy as pseudo-labels. Figure 2 shows their comparison: in Family Relationships and Bitwise Arithmetic, we see larger gains in majority voting performance, and likewise SRT outperforms its fixed teacher variant substantially, by 10% for Bitwise Arighmatics, 8% in Family Relationship, and 6% in Knights and Knaves. On Knights & Knaves, the starting policy already has >90% majority voting accuracy, and we see the difference between the evolving teacher and fixed teacher variants of SRT to be smaller here. Furthermore, these performance gains cannot be explained by learning to format properly: since the model was trained on an easier difficulty level, it can already format its responses correctly in most cases (Appendix E), and we see performance climbing after saturation in model’s format correctness. Overall, on the synthetic reasoning tasks, SRT clearly pushes the model beyond its starting capabilities, showing the promise of self-improvement even from this basic recipe for self-supervision.
We show a summary of our empirical findings using different base models and training datasets in Figures 3 and 4 (for training curves using more combinations of (train, test) dataset pairs, refer to Appendix C.1 and C.2). On all instances, SRT improves both avg@16 and majority vote@16 accuracies on heldout MATH-500 prompts, and performs on par with regular RL training with ground truth verification. More impressively, the observations hold for base models like Llama-3.1-8B-Instruct, which is known to be particularly difficult for RL training on reasoning tasks (Gandhi et al., 2025), improving its average accuracy from 52.6% to nearly 60%.
We also compare SRT with its offline variants: SFT on the majority vote (Huang et al., 2023), DPO and ScPO (Prasad et al., 2024) employing contrastive learning between the majority vote and non-majority vote answer in Table 1. We observe that SRT retains better performance compared to its offline variants distilling the majority vote decisions of the base policy, showing the benefit of self-improvement in the label-generating policy. For more details about the baselines, including choice of hyperparameters and performance on additional test datasets, refer to Appendix F.
Table 1: (Peak Performance) We compare SRT against various baselines using a Qwen2.5-Math-7B model. Reported numbers are peak mean@32 accuracy, averaged over 3 heldout datasets: AIME 2024, AIME 2025, AMC. We find that SRT generally retains better performance compared to baseline methods distilling the majority voting decisions of the fixed teacher policy into the current model.
Train Data Base Model SFT DPO ScPO SRT RL on Ground Truth
MATH-12K 0.15 0.18 0.23 0.20 0.32 0.33
DAPO 0.15 0.18 0.21 0.20 0.31 0.36
Takeaway 1: SRT can improve reasoning capabilities beyond the base model On both synthetic and real reasoning tasks, SRT improves average and majority voting accuracies, showing ability gains beyond the base model. Specially, improvement in majority voting accuracy also signifies improvement in the quality of generated self-supervision during training, demonstrating the importance of training the “verifier” in addition to just training the policy. Overall, our simple recipe shows a promising path forward to self-improvement.
4.2 Can Self-Improvement from SRT be Sustained Indefinitely?
Given the strong performance of SRT in various reasoning tasks and model architectures, an important question is whether self-training can be maintained over extended iterations. Similar to the prior section, we first test SRT on synthetic tasks with controllable difficulty to rigorously study its properties, and then test the resulting insights on real-world reasoning domains.
Figure 5: (Multi-level climbing on Reasoning Gym using curriculum) The Qwen3-4B-Base model can climb on progressively more difficult tasks without ground truth labels via a simple curriculum strategy — where we train an earlier level’s final checkpoint with SRT on the next difficulty level. This approach also seems to improve both average and majority voting accuracy on each level.
Multi-level self-improvement on synthetic tasks using curriculum.
Since Reasoning Gym provides a built-in way to control the difficulty of generated tasks, we first investigate whether self-training on an easier set of tasks can produce a model capable of self-improvement on progressively harder levels of difficulty. To do so, we choose 2 Reasoning Gym tasks: Bitwise Arithmetic and Knights & Knaves. Similar to our previous setting, we first train using RL on ground truth labels on the easiest level of difficulty, then progressively train on harder levels without ground truth labels (i.e., SRT on level 5 starts with the checkpoint obtained from SRT on level 4, and so on). For more details, refer to Appendix B.3.
Figure 5 shows our primary results: in this controlled setting, SRT is able to maintain self-improvement on progressively harder difficulties. In particular, SRT can show reasonable improvement in Bitwise Arithmetic Level 4 after being initialized on Level 2 with ground-truth training, and also progressively climb to near 100% accuracy on Knights & Knave Level 9 after being trained with ground-truth on Level 2 only (intermediate levels are trained with SRT). In synthetic reasoning tasks where we have a knob for controlling the difficulty of training problems, even the simple curriculum of progressively training on more difficult problems with self-supervision can maintain improvement across multiple difficulty levels.
Figure 6: (Extended self-training leads to model collapse) Inspired by multi-level improvement on reasoning gym tasks, we take four LLMs with strong math abilities from pretraining, and train them with SRT for an extended period of time. SRT improves performance at first, but then demonstrates complete model collapse on all 4 base models. (Note: on Llama-3.1-8B-Instruct, the learning rate used in Figures 3 and  4, $10^{-7}$, does not lead to model collapse within our training budget, but $3\times 10^{-7}$, a slightly higher learning rate does — we hypothesize that with an extended training run, even $10^{-7}$ would lead to model collapse. For more details on the effect of hyperparameters on model collapse, refer to Figure 8.)
Extended SRT-training on math problems.
Next, we test whether our insights from synthetic tasks also extend to real-world math problems. Specifically, we take take the same 4 base models as in the previous section, and train them on a difficult math dataset through an extended number of iterations using SRT. Figure 6 demonstrates our surprising finding: while SRT initially increases the base model’s performance at a comparable rate with ground-truth RL training, extended training using SRT leads to sudden performance collapse. We observe performance collapse or degradation from extended SRT-training across all models and training datasets, and record these in detail in Appendix C.1 and C.2. Given that this surprising phenomenon deserves more investigation, we study it in more detail next.
What happens after SRT peaks in performance?
To develop a clearer understanding of the underlying reasons for this phenomenon, in this section we investigate this SRT-induced model collapse closely.
We plot the training statistics of the SRT objective (Eqn. 6) in Figure 7. The observed performance collapse closely coincides with a sudden increase in the SRT self-reward objective, implying that the optimization procedure has, in fact, maximally optimized the training objective (self-consistency majority voting), despite a decline in actual output correctness. On the same figure, we further report the token-level average Kullback–Leibler divergence between the model under SRT training and the base model. We observe a sharp increase in KL divergence at the exact point when performance begins to deteriorate, indicating that the generative distribution of the model has substantially diverged from the original model.